what is the limit of 3x^2+3ax-2a^2 over x^2-a^2 as x approaches to a.
1
Expert's answer
2020-06-08T20:15:36-0400

The limit is given by:

"\lim_{x \to a} \frac{3x^2 + 3ax - 2a^2}{x^2 - a^2}"

"\\"

"\\" Left Hand Limit:

"\lim_{x \to a^{-} } \frac{3x^2 + 3ax - 2a^2}{x^2 - a^2}"

From the function of limit and the limit value, we can say that the function decreases without a bound. So, we can write:

"\lim_{x \to a^{-} } \frac{3x^2 + 3ax - 2a^2}{x^2 - a^2} = - \infty"

"\\"

Right Hand Limit:

"\lim_{x \to a^{+} } \frac{3x^2 + 3ax - 2a^2}{x^2 - a^2}"

From the function of limit and the limit value, we can say that the function grows without a bound. So, we can write:

"\lim_{x \to a^{+} } \frac{3x^2 + 3ax - 2a^2}{x^2 - a^2} = \infty"

"\\"

Since:

"\lim_{x \to a^{-} } \frac{3x^2 + 3ax - 2a^2}{x^2 - a^2} \neq \lim_{x \to a^{+} } \frac{3x^2 + 3ax - 2a^2}{x^2 - a^2}"

Hence:

"\lim_{x \to a} \frac{3x^2 + 3ax - 2a^2}{x^2 - a^2}" does not exist.